### v2

Abstract
We explore the application of machine learning algorithms to predict the suitability of Russet potato clones for advancement in breeding trials. Leveraging data from manually collected trials in the state of Oregon, we investigate the potential of a wide variety of state-of-the-art binary classification models. We conduct a comprehensive analysis of the dataset that includes preprocessing, feature engineering, and imputation to address missing values. We focus on several key metrics such as accuracy, F1-score, and Matthews correlation coefficient (MCC) for model evaluation. The top-performing models, namely the multi-layer perceptron (MLP), histogram-based gradient boosting (HGB), and support vector machine (SVM) classifiers demonstrate consistent and significant results. Variable selection further enhances model performance and identifies influential features in predicting trial outcomes. The findings emphasize the potential of machine learning in streamlining the selection process for potato varieties, offering benefits such as increased efficiency, substantial cost savings, and judicious resource utilization. Our study contributes insights into precision agriculture and showcases the relevance of advanced technologies for informed decision-making in breeding programs.
keywords:
Russet potatoes, machine learning, classification, variable selection
1 Introduction
Potato (Solanum tuberosum L.) is an important food crop in the United Sates with an annual potato production of 392.24 million cwt in 2022. The Pacific Northwest forms the largest sub-region in potato production in the U.S., contributing 60.1% of U.S. potato production according to USDA’s National Agricultural Statistics Service. Potatoes are not only consumed as a fresh vegetable but also form a raw material for the French fry, chip, and starch processing industries. In the U.S., 39% of the potato crop is utilized by the frozen processing sector. The Columbia basin of Oregon and Washington states is the leader in the frozen potato sector, with more than 80% of potatoes grown here being russet type for utilization by the processing industry.
Developing new russet potato varieties with high yield, good processing quality, climate resilience, and disease resistance is the main focus of the Oregon State University’s Potato Breeding and Variety Development program, an integral part of the Northwest Variety Development Program (also known as the Tri-State program). Typically, it takes 12–13 years from the hybridizations to select and release a new potato variety. In the first field year, seedling tubers generated from true potato seeds are planted in Klamath Falls (around 60,000 clones) as single hills and about 1% of the clones selected are planted at Hermiston and Klamath Falls in the second field year. Phenotypic field selections are performed in the first and second years and then the selected clones are tested through Oregon statewide replicated trials at Hermiston, Klamath Falls, and Ontario. The Oregon statewide trials are typically performed in years three to five and the selected clones will then move to a Tri-State trial, and then to regional trials from the sixth or seventh field year to the tenth or eleventh year. Potato varieties released from the Tri-State variety development program such as the “Clearwater Russet”, “Mountain Gem Russet”, “Umatilla Russet”, and “Ranger Russet” have proved to be significantly successful by producing higher total yield and higher percentage of finished product derived from raw tubers in the processing industry. Benefits from the adoption of new russet potato varieties continue to grow over time [1].
In the breeding program, the selected potato clones must be grown and evaluated in research replicated trials for multiple years. Each selected clone will be planted in diverse locations and agronomic data is collected and evaluated for its performance. Data for each potato clone collected annually includes field data such as emergence, plant vigor, plant uniformity, stems per plant, vine maturity; after harvest data includes total yield, yield for different weight categories; attributes pertaining to the potato’s external characteristics, specific gravity, internal defects, and performance following the process of frying. The extensive costs associated with managing these selection trials and the subsequent data collection after harvest account for a significant proportion of investment in the Oregon Potato Variety Development Program.
The complexity further increases due to the intricate interplay between distinct traits and environmental interactions needed for multi year and multi location trials. Breeders must decide whether the potato clone should go into the subsequent selection trial stage based on its performance or whether retaining a poorly performing clone could add unnecessary cost to the program resulting in decreased breeding selection efficiency. Advances in genetic tools have resulted in the use of molecular markers and genomic selection to improve selection efficiency. Even with molecular and genomic data, it is essential to carry multi year, multi location trials and the breeder must select which clones to advance. In this context, artificial intelligence (AI) or machine learning technologies can help evaluate potato clones and assist in making such decisions [7]. The ascendancy of AI technologies in revolutionizing data analytics is evident, particularly within breeding programs. Researchers have used machine learning models to obtain highly accurate predictions of root yield based on the vegetative indices of the different growth stages of cassava (Manihot esculenta Crantz) [18]. With the potent integration of machine learning algorithms and chemical composition as predictors, researchers have successfully predicted consumer flavor perceptions of blueberries and tomatoes [4]. Data-informed decision making facilitated by machine learning models can assist breeders select potato clones more accurately and reduce the number of varieties proceeding to subsequent stages. The consequential impact includes substantial financial savings, increased breeding selection efficiency, and judicious resource utilization.
1.1 Related Work
Precision agriculture and the incorporation of advanced technologies like artificial intelligence are becoming crucial for maintaining high-quality yields for many crops. Machine learning methods have been applied to many crops whose information was obtained through various data collection techniques [6, 2].
In a study of French fry potato in Poland, machine learning methods of regression to predict yield using diverse datasets (agronomic, weather, satellite, and soil data) were used [11]. The authors found that consideration of features, coupled with appropriate data preprocessing techniques (e.g., imputation) and outlier management, is crucial for accurate predictions in the field of agricultural yield forecasting.
Another study of potato employed the machine learning models of support vector machines (SVM) and multi-layer perceptrons (MLP) to classify potato varieties based on macroscopic observations and measurements for selection or de-selection during trials much like those used in our study [13]. For each of these two classification techniques, four models were constructed—two for binary classification and two for 4-class classification. The results were reported in terms of the accuracy score (number of correct predictions) with the SVM method demonstrating higher accuracy than MLP in the 4-class models. It is unclear how many features [13] considered or how they addressed missing data. Further, no parameter tuning was performed.
2 Methods
Code for the methods described here can be found in our repository: https://github.com/fabstat/burbank.
2.1 Study Area and Data
From 2013–2014 there were four regions in Oregon, represented in Figure 1, in which the Russet potato trials were conducted: Hermiston (HER), Ontario (ONT), Klamath Falls (KF) and Corvallis (COR). Beginning in 2015, COR was dropped from the russet trials and two trials were conducted in HER, early and late (see Table 1).
Figure 1: US State of Oregon. Points represent the cities in which the potato trials were conducted: Corvallis, Hermiston, Ontario, and Klamath Falls. From each of these states, russet potatoes such as the one depicted at the center of the map (source: Wikimedia) were collected and data of each clone’s characteristics were recorded for analysis.
Table 1: Early and late planting season dates over the years.
Early Planting Late Planting
Year Planting Vine Kill Planting Vine Kill
2013 04/05/13 09/04/13
2014 04/16/14 09/11/14
2015 03/26/15 07/22/15 04/06/15 09/04/15
2016 03/25/16 07/29/16 04/07/16 09/02/16
2017 03/27/17 08/10/17 04/04/17 08/28/17
2018 03/26/18 08/08/18 04/09/18 08/31/18
2019 03/28/19 08/01/19 04/16/19 09/02/19
2020 03/23/20 08/01/20 04/21/20 08/31/20
2021 03/17/21 07/30/21 03/31/21 08/31/21
[…]
The preprocessed dataset includes 1086 total clones, 216 of which are unique. Each clone is trialed 1–3 years. Most clones were in their first year of trial. There are three control varieties planted in all regions and all years: Ranger Russet, Russet Burbank, and Russet Norkotah. Two other control varieties are tested only in Hermiston: Shepody (early) and Umatilla (late), and were measured for some but not all years. For each clone, data for 38 attributes are recorded in a spreadsheet for each particular region. This data pertained to the total yield, yield per size category, features related to the potato’s appearance, and features related to the potato after being fried. See Table A for more details on each feature.
2.2 Feature Engineering
Not all 38 aforementioned features were used in our analysis, and some additional features were included. Features with over 400 not applicable (NA) ratings were dropped: “flower color”, “vine size”, and “maturity”. The feature “bruise data” was also dropped as it contained mostly text description and was missing for most observations. The following new features were added:
  1. 1.
    “true_keeps”: response variable. A clone is assigned a “1” if it passed three trials and a “0” if it passed fewer than three trials. Clones in the first or second year of the statewide trials during this analysis were removed from the dataset.
  2. 2.
    “Ctrl ave”: Average total yield of control clones for a particular year and trial region. All clones from a particular year and trial region should have the same “control average”.
  3. 3.
    “CA”: 100×(Total YieldCtrl $100\times\left(\frac{\textnormal{Total Yield}}{\textnormal{Ctrl Ave}}\right)$ represents the percentage of a particular variety’s total yield from the averaged controls’ total yield.
  4. 4.
    “GDD 1–60”: the sum of the first 60 growing degree days
  5. 5.
    “GDD 61–90”: the sum of growing degree days from day 61 to day 90
  6. 6.
    “GDD 91–end”: the sum of growing degree days from day 91 to vine kill date
The last three variables pertain to growing degree days (GDD) defined as
$\mathrm{GDD}=\left[\frac{(T_{\mathrm{MAX}}-T_{\mathrm{MIN}})}{2}\right]-T_{\mathrm{BASE}}$
where $T_{\mathrm{MAX}}$ and $T_{\mathrm{MIN}}$ are the day’s maximum and minimum air temperatures, respectively, and $T_{\mathrm{BASE}}$ is “the temperature below which the process of interest does not progress” [12]. In our case, $T_{\mathrm{BASE}}$ is the lowest temperature at which potato continues to grow, or 8 Celsius.
2.3 Data Imputation
While the overall dataset was reasonably large, a significant number of variables were missing values (or entries) for several observations. For instance, the variables “flower color” and “maturity” were missing for 638 and 639 observations, respectively (of the total of 1086 clones). When considering only observations with numeric yield, 25 of the 36 original attributes have at least 1 missing missing value, with 13 of these features each having more than 100 missing values. Hence, imputation was necessary to consider the complete dataset. Without imputation, one must drop observations with missing values. Since a few of the attributes were not recorded for almost half of the dataset, dropping observations after data processing and cleaning would reduce the set from 1086 observation to 209 observations. Such a small set of observations would lead to overfitting and machine learning that might not be meaningful.
To impute data, we first standardized the dataset and then used Scikit-learn’s [14] function IterativeImputer with sample_posterior=True, which performs multiple imputation by chained equations, or MICE [20]. Under the assumption that the data is missing at random, MICE works by first performing a simple mean imputation to serve as place holders for all variables. For each variable in turn, the place holder means are removed and the variable is treated as dependent on the other independent variables in a regression problem.
Besides the imputed dataset, we also constructed a non-imputed dataset where we chose a maximum number of “NA’s” any given column could have so that the dataset was reduced to at most the largest number of NA’s that was below this maximum threshold. In the analysis presented below, an NA threshold of 400 was selected to ensure we kept at least two thirds of the dataset even when dropping rows containing NA’s. For purposes of comparison, we constructed both the imputed and non-imputed datasets by first dropping columns with over 400 NA’s. Postprocessing there were 885 observations in the imputed set and 404 observations in the non-imputed set, with the same set of 40 features in each.
Once numerical variables were standardized, we hot encoded the categorical variables - a process where all categorical variables’ values are accounted for all observations, but for an observation’s given category, its value is encoded to 1 while the other non-applicable values are encoded to zero - and handled missing values in a similar manner, i.e., they were either dropped or imputed. We then ran a grid search with cross validation to find the best set of hyperparameters for each of the models considered (see Section 2.5).
2.4 Modeling Approach
With the ultimate goal of automating the process of discriminating varieties that meet or exceed expected targets from those that do not meet the standards specified above, we explored a large family of machine learning classification models on our dataset and drew structural insights from their results. Note that this process is currently fully manual, where the potato breeding expert looks at every datum in order to make a decision. As such, having a more automated way to make, or aid in making, the decision is of great advantage. We concentrated on nonlinear classifiers since we expected them to perform as well or better than linear ones. Nonlinear classifiers are more flexible. They do not rely on a line, or hyperplane in higher dimensions, to separate the data classes, and they are more robust against outliers.
Clones must be selected for breeding in all three years of the Oregon statewide trials, if they are to graduate to the Tri-state trials. We labeled each clone in the dataset as “1” if it passed three trials and as “0” if it passed fewer trials. Clones were removed from the dataset if still in the first or second year of statewide trialing during this analysis. We performed model selection and hyperparameter tuning via grid searches using 5-fold cross-validation and split the data such that 80% was used in training and 20% was used for testing.
2.5 Classification Models and Evaluation Metrics
Since the goal was to separate two classes in a high dimensional feature space efficiently, we constructed a large family of nonlinear classification models. As these classification methods may not be widely used for agricultural data analysis, we present high level explanations for each of them below.
  1. 1.
    $\boldsymbol{k}$-Nearest neighbors: This is a fundamental and widely applied clustering method in unsupervised learning. The objective of the algorithm is to partition a dataset into $k$ clusters (for a predefined value of $k$) based on similarity, by iteratively computing the Euclidean distance of each data point and cluster representatives, assigning the data points to clusters, and then moving the cluster representative to be the center of the data points it represents. Such an algorithm leads to a local optimum and so it is sensitive to initial conditions and shape assumptions. Also, since clustering relies on computing arithmetic means, this algorithm is sensitive to outliers [17].
  2. 2.
    Gaussian process based on Laplace approximation: This approach involves placing a Gaussian Process (GP) prior over a latent function, which is then transformed through the logistic function to obtain a prior on the class probability. The Laplace approximation is employed for posterior inference, approximating the non-Gaussian joint posterior with a Gaussian distribution. With this method, positive training examples will give rise to a positive coefficient for their kernel function, while negative examples will give rise to a negative coefficient, making it analogous to the support vector machine (SVM) model (see Method 10 below). However, all features provide information to perform predictions, making it an inefficient algorithm in high dimensional spaces [21].
  3. 3.
    Gaussian naïve Bayes: This is a naïve Bayes method where the likelihood of the features is assumed to be Gaussian. A naïve Bayes method is a supervised learning technique grounded in the application of Bayes’ theorem, which incorporates the “naïve” assumption of conditional independence among all pairs of features given the class variable $y$ and features $x_{1},\dots,x_{n}$ as follows:
    $P(y|x_{1},\dots,x_{n})=\frac{P(y)P(x_{1},\dots,x_{n}|y)}{P(x_{1},\dots,x_{n})}$
    $\Rightarrow$ $P(y|x_{1},\dots,x_{n})\propto P(y)\prod\limits_{i=1}^{n}P(x_{i}|y).$
    The predicted class variable maximizes the right-hand side of the proportion, given the relative frequency of the class variable in the training dataset [22].
  4. 4.
    Quadratic discriminant analysis: This supervised learning technique computes discriminant functions for each class, which are then used to determine the decision boundary between each class. Assuming each class density has a multivariate Gaussian distribution and by comparing the log-ratio of two classes posteriors, we can obtain the quadratic discriminant functions
    $\delta_{k}(x)=-\frac{1}{2}\log|\boldsymbol{\Sigma}_{k}|-\frac{1}{2}(x-\mu_{k})^{T}\boldsymbol{\Sigma}_{k}^{-1}(x-\mu_{k})+\log\pi_{k}\,.$
    The decision boundary between each pair of classes $k$ and $l$ is the set of $x$ values satisfying the quadratic equation $\delta_{k}(x)=\delta_{l}(x)$ [9].
  5. 5.
    Decision tree: This is a supervised learning approach in which a given vector of explanatory variables $X$ is partitioned into the nodes of a binary tree whose edges represent the mutually exclusive set of outcomes of the parent node. Each edge leads to a different input feature until it reaches the leaf nodes, where a class is assigned to the sequence of feature node outcomes; the goal is to group samples with the same label. Classification depends on the underlying decision tree structure, which is decided by hyperparameters such as the split criteria and the maximum depth of the tree. Although simple to understand, decision trees can be unstable or cause over-fitting depending on the choice of hyperparameters [9].
  6. 6.
    Adaptive boosting: This is a commonly used ensemble learning approach in which a sequence of weak learners (models that are slightly better than guessing randomly), such as small decision trees, is repeatedly applied to modified versions of the data (i.e., by changing the weights applied to the training observations). The predictions from all the weak classifiers are then combined to produce a majority vote prediction [9].
  7. 7.
    Random forest: This ensemble learning approach is an extension of the bootstrap aggregating (or bagging) method, in which random samples of the training data are selected with replacement and then trained on individually using the base learner. Random forests work similarly by constructing an uncorrelated collection of decision trees. Each tree in the ensemble is constructed from training data sampled with replacement, and an exhaustive search through the features is performed in order to obtain the best split in the node set. Although this is a robust algorithm and might be selected for this reason, it is more complex and time consuming than other ensemble techniques [3].
  8. 8.
    Multi-layer perceptron (MLP): This is a feedforward neural network with at least three layers—an input layer, one or more hidden layers, and an output layer—in which neurons are fully connected with a nonlinear activation function. The MLP classifier is trained by backpropagation to optimize a loss function. The Broyden–Fletcher–Goldfarb–Shanno (BFGS) method, the optimization algorithm used in this framework, is an iterative method for solving unconstrained nonlinear optimization problems [8].
  9. 9.
    Histogram-based gradient boosting (HGB): Gradient boosting algorithms are additive machine learning models that combine many weak learners such as decision trees to form a strong learner. As finding the best split points for learning the decision trees via sorting the values of each feature is time and memory intensive, the histogram based algorithm buckets continuous feature values into discrete bins and constructs a histogram, thus reducing the number of unique values to sort [10]. Several hyperparameters in the histogram-based gradient boosting classifier (HGBC) can be optimized during learning, such as maximum tree depths, maximum number of bins and the $L_{2}$ regularization parameter, to avoid over-fitting issues while affording good performance.
  10. 10.
    Support vector machine (SVM) with a radial basis function kernel: A linear support vector machine (SVM) identifies the plane that maximizes the distance between itself and the closest set of points (specifically the feature vectors from a set of observations) that belong to the two distinct categories. Figure 2(a) shows a linearly separable dataset with a linear SVM decision boundary.
    (a) Two classes of points separated by a line.
    (b) Two classes of points separated by a circular curve.
    Figure 2: A dataset where the two classes of points are separable with a line (Left), and one where a nonlinear kernel (radial basis function or RBF in this case) does the separation (Right) when no linear separation is possible.
    If the data is not linearly separable for the desired classification problem, then there will be no hyperplane separating the two classes. However, a nonlinear decision boundary may still exist. As SVM is an inherently linear classification method, we transform the vector space via a kernel trick. A kernel is a function that take as inputs the observations of the original feature space and outputs the dot products of vectors in a transformed target space where finding hyperplanes that separate the data is more likely [16]. Two important hyperparameters in a nonlinear SVM are $C$ and $\gamma$. The scale parameter $\gamma$ determines the radius of influence of a single observation, while $C$ controls the size of slack variables that allow some observations to be outside the target hyperspace. Setting these parameters appropriately is a trade-off between controlling the classification error and overfitting the data.
  11. 11.
    An ensemble of classifiers along with a meta-classifier: Also called a stacking classifier, two or more classifiers are applied to the datasets and the resulting classification predictions are used to train a meta-classifier [19]. The stacking classifier used in this study has gradient boosting and SVM as its level-1 classifiers, and logistic regression as its level-2 (or meta-) classifier.
2.5.1 Evaluation Metrics
To evaluate the quality and performance of the large family of classification models, we computed several metrics including accuracy (proportion of correct predictions), F1-score (harmonic mean of precision and recall), and Matthews correlation coefficient (MCC), which is defined as:
MCC $=\frac{\text{TP}\times\text{TN}~{}-~{}\text{FP}\times\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}\,,$ (1)
where TP is the number of true positives, TN the number of true negatives, FP the number of false positives, and FN the number of false negatives. The MCC ranges from $-1$ to $1$ and on imbalanced datasets, it can be more informative than either accuracy or the F1-score [5]. Our imbalanced dataset includes 88 positive examples and 998 negative examples.
2.5.2 Variable Selection
We performed variable selection on both the imputed and non-imputed datasets for the top-performing classification models used in this study. For each model described in Section 3.1, we performed a grid search with cross-validation on the hyperparameter space to determine the best scoring model. We then performed forward feature selection by beginning with zero features in our models and iterating through each available feature, computing each model’s score with that feature included in it. The feature resulting in the highest model score becomes a feature in the model, and the search for a second feature proceeds. The process continues until we have the desired number of features in the model or an additional feature does not significantly improve the model score. The results from variable selection are presented in Section 3.1.
3 Results
Table 2 presents the MCCs for the classification models explored in this study on the imputed and non-imputed datasets. Columns containing more than 400 NA’s were dropped before imputation and data cleaning, so that both the imputed and non-imputed datasets contained the same set of features. The top performing classifiers (highlighted in gray) were the multi-layer perceptron (MLPC), histogram-based gradient boosting (HGBC), and support vector machine (SVC), as each obtained an MCC score above or very close to 0.5 in the cross-validation and test for both imputed and non-imputed datasets with high significance. We note that AdaBoosting achieved similar performance to the declared top performing models, however we will not discuss it further as we will focus on HGBC, which uses the same underlying principle of an ensemble of weak learners but it is a more modern algorithm allowing additional hyperparameters to be set (e.g., regularization).
Models whose test MCC is very close to zero are no better than a random guess. Gaussian process, Naïve Bayes, QDA, and Decision tree classifiers all performed particularly poorly on our datasets. Although Gaussian process is analogous to SVM, its implementation used in this analysis is non-sparse, and hence the model uses all features for each example when computing the high-dimensional posterior. As such, a possible explanation for this model’s poor performance is that the posterior computed from the training examples is skewed. If the posterior is skewed, the Laplace approximation (which is centered on the mode) will not be accurate [15].
The poor performance from Naïve Bayes and QDA may be due to the assumption of feature independence inherent to these models. An inspection of the table of features (Table A) makes it clear that several variables considered are not independent. Lastly, a decision tree’s performance is very sensitive to its structure, where features are greedily made into nodes. It is possible that variable selection and a more thorough search of the hyperparameter space could allow these models to better perform on our dataset. However, here we focus on those models showing better promise under the same circumstances.
Table 2: Best MCCs for all the classification models considered in this study. Top performing models (highlighted in gray) were multi-layer perceptron (MLPC), histogram-based gradient boosting (HGBC) and support vector machine (SVC), on datasets with both multiple imputation and without imputation. The results presented here are from models using all available features which have less than 400 NA’s.
Imputed Not Imputed
Model Test MCC CI Test MCC Test MCC CI Test MCC
$k$-NN 0.413 (0.378, 0.447) 0.338 (0.283, 0.392)
Gaussian process 0.000 (--0.042, 0.042) 0.000 (--0.062, 0.062)
Naïve Bayes 0.373 (0.336, 0.408) --0.004 (--0.066, 0.058)
QDA 0.288 (0.250, 0.326) 0.000 (--0.062, 0.062)
Decision tree --0.035 (--0.077, 0.006) 0.089 (0.027, 0.150)
AdaBoosting 0.523 (0.492, 0.553) 0.541 (0.496, 0.583)
Random forest 0.085 (0.043, 0.126) 0.531 (0.486, 0.574)
MLPC 0.530 (0.500, 0.559) 0.569 (0.525, 0.609)
HGBC 0.563 (0.534, 0.591) 0.485 (0.436, 0.531)
SVM 0.563 (0.534, 0.591) 0.692 (0.658, 0.723)
Stack 0.534 (0.504, 0.563) 0.432 (0.380, 0.481)
[…]
3.1 Top Performing Models
From among the large family of classification models considered (Section 2.5), three models performed consistently well with both imputed and and non-imputed data. They were the multi-layer perceptron classifier (MLPC), the histogram-based gradient boosting classifier (HGBC), and the support vector machine classifier (SVC). The results from these models without variable selection are highlighted in Table 2 in light gray, and their confusion matrices are presented in Figure 3. We explored forward feature selection on these three models to understand which variables are important in determining whether a potato clone is kept in the trials. The resulting MCC scores for the top three performing models using forward feature selection on the imputed and non-imputed datasets are presented in Table 3.
Figure 3: Confusion matrices for the top performing models applied to both imputed and non-imputed datasets: multi-layer perceptron classifier (MLPC, called a Neural Net here) in first row, histogram-based gradient boosting classifier (HGBC) in middle row, and support vector machine (SVM) classifier in the last row. Variable selection was not performed here, however variables containing at least 400 NA’s were dropped before training. Numerical values within the confusion matrices represent prediction proportions.
Table 3: MCC for the top classification models after forward feature selection.
Imputed Not Imputed
Model $\Delta$CV MCC Test MCC CI Test MCC $\Delta$CV MCC Test MCC CI Test MCC
MLPC 0.476 0.608 (0.581, 0.633) 0.595 0.623 (0.584, 0.659)
HGBC 0.531 0.574 (0.546, 0.602) 0.511 0.645 (0.608, 0.680)
SVM 0.539 0.502 (0.471, 0.533) 0.731 0.623 (0.584, 0.659)
[…]
It is interesting to note that variable selection significantly improved the test MCC of MLPC on both imputed and non-imputed datasets, as well as the test MCC of HGBC on the non-imputed dataset. However, this form of variable selection seems to hinder the performance of the SVM classifier on the imputed and non-imputed datasets. One reason for this drop in performance could be that features were selected after the hyperparameter search. Both techniques use the SVM classifier, which in turn suppresses certain features through an $\ell_{2}$-norm penalty. Thus, the second application of SVM for feature selection would drop the variable with high penalty regardless of their overall contribution to the decision function.
Table 4 lists the variables that were selected for each of the three top-performing models on each imputed and non-imputed dataset. We highlight in bold the variables selected in all three models when applied to a particular dataset. On the imputed dataset the variables “Trial Region_COR”, “Trial Region_HER_Early”, “Yr in trial”, “%RB”, “Vascular”, “GDD 61–90” and “GDD 91–end” were selected for all three models. On the other hand, the common variables selected for the non-imputed dataset were “Trial Region_HER_Late”, “Yr in trial”, “Hollow Heart”, “Brown Center”, ‘Vascular”, “Shape Uniformity”, and “GDD 1–60” instead.
Two variables selected in the models using the imputed dataset, “Trial Region_COR” and “Trial Region_HER”, gained their significance due to only being present for the years 2013 and 2014. Respectively, they accounted for 4.9% and 5.2% of the total kept varieties for all regions and years present in the imputed dataset. This imbalance may be the sole reason these variables were selected as informative.
The variables selected by all three models and for both datasets were “Yr in Trial” and “Vascular”. The importance of “Yr in Trial” may be irrelevant as it indicates how long a particular potato variety survived selection during the three-year trial. “Vascular” refers to vascular discoloration, a darkening of the ring of tissue that surrounds the tuber just below its skin. The ring contains the tuber’s vascular system, and its discoloration can be indicative of bacterial or fungal infections or abiotic stress response. The vascular discoloration in tubers can lead to low processing quality and is an important trait for selection.
Table 4: Variables selected via forward feature selection (second column) using the corresponding top performing algorithm on each dataset (first column). Bold variable names highlight the variables that were selected in common between the algorithms when applied to the same dataset (imputed versus not imputed).
Model Variables Selected
MLPC (imputed dataset) “Trial Region_COR”, “Trial Region_HER_Early”, “Trial Region_ONT”, “Yr in trial”, “Total Yield”, “%RB”, “%RB 1s”, “%1s”, “4-6oz”, “Y2s + Over 20oz”, “Culls”, “Specific Gravity”, “Blackspot”, “IBS”, “Vascular”, “Tuber Shape”, “Eye Depth”, “Ctrl Ave”, “CA”, “GDD 1–60”, “GDD 61–90”, “GDD 91–end”
HGBC (imputed dataset) “Trial Region_COR”, “Trial Region_HER”, “Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_KF”, “Trial Region_ONT”, “Yr in trial”, “%RB”, “Over 20oz”, “4-6oz”, “Under 4oz”, “Culls”, “Tuber/Plant”, “Sugar Ends”, “Hollow Heart”, “Brown Center”, “IBS”, “Vascular”, “Russeting”, “Ctrl Ave”, “GDD 61–90”, “GDD 91–end”
SVM (imputed dataset) “Trial Region_COR”, “Trial Region_HER”, “Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_KF”, “Trial Region_ONT”, “Yr in trial”, “Total Yield”, “%RB’’, “Y1s”, “%RB 1s”, “%1s”, “Y2s + Over 20oz”, “Specific Gravity”, “Fry Color Stem”, “Fry Color Bud”, “Sugar Ends”, “Blackspot”, “Vascular”, “Scab”, “GDD 61–90”, “GDD 91–end”
MLPC (non-imputed dataset) ‘Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_ONT”, “Yr in trial”, “%RB 1s”, “10-20oz”, “6-10oz”, “4-6oz”, “Length/Width”, “Specific Gravity”, “Fry Color Stem”, “Hollow Heart”, “Brown Center”, “Vascular”, “Tuber Shape”, “Shape Uniformity”, “Growth Cracks”, “Ctrl Ave”, “GDD 1–60”, “GDD 61–90”, “GDD 91–end”
HGBC (non-imputed dataset) “Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_ONT”, “Yr in trial”, “Total Yield”, “%RB”, “Y1s”, “%RB 1s”, “Over 20oz”, “10-20oz”, “Culls”, “Tuber/Plant”, “Hollow Heart”, “Brown Center”, “Blackspot”, “Vascular”, “Shape Uniformity”, “Eye Depth”, “Scab”, “Shatter Bruise”, “GDD 1–60”
SVM (non-imputed dataset) “Trial Region_HER_Late”, “Yr in trial”, “Total Yield”, “%RB”, “Y1s”, “%RB 1s”, “%1s”, “Tuber/Plant”, “Length/Width”, “Specific Gravity”, “Hollow Heart”, “Brown Center”, “Vascular”, “Skin Color”, “Tuber Shape”, “Shape Uniformity”, “Scab”, “Ctrl Ave”, “GDD 1–60”, “GDD 61–90”, “GDD 91–end”
[…]
4 Discussion
We have applied a large collection of machine learning classification techniques to manually collected Russet potato trial datasets. The main goal was to investigate whether and how machine learning methods might provide domain experts with faster and more efficient ways to identify potato varieties for advancement to later trials. We discovered that although missing data can pose a challenge, we can use imputation or strategic dropping of features, as well as engineer new features, to obtain significant improvements in the model classification performances. As illustrated in the results, the large number of negative examples in our dataset resulted in strong performance for predicting when to drop varieties from the trials. Although this bias in the predictions can lead to false negatives, it also means that varieties that will potentially fail in future trials are dropped earlier on, making the process more efficient and economical. Further, our results suggest that integrating machine learning techniques with the art of selection can result in increased selection efficiency. As the variety trials are costly and keeping a clone that is not worthy for one additional year can be cost detrimental. Our results indicate, with availability of extensive data, an integrated art of selection along with data-informed confirmation from machine learning can be effective.
In conclusion, this work has demonstrated the potential of machine learning methods in revolutionizing potato variety selection during trials. The integration of varied data sources, coupled with advanced data-processing techniques, present a pathway toward a more efficient and informed selection process.
Appendix A Preprocessed Dataset Characteristics
Table A: Attributes for which data was collected and averaged across 4 replicates of each unique clone.
Attribute Name Attribute Description
Yield Total in hundredweight/acre. Sum of all different sizes, yield 2’s and culls where each has been multiplied by 8.31. Dataset ranges from 0-1510 with mean of 618. There are 47 missing values.
Rank An integer from 1 to number of varieties for a specific location and year. It ranks from highest (1) to lowest total yield. Ranges from 1-44. There are 47 missing values.
% RB Total yield divided by Russet Burbank’s total yield multiplied by 100. Ranges from 0-274. There are 47 missing values.
Yield No. 1’s Sum of all yields (in cwt) from which the potatoes are between 4 and 20 ounces. Ranges from 0-983. There are 47 missing values.
Rank (No. 1’s) An integer from 1 to number of varieties for a specific location and year. It ranks from highest (1) to lowest yield No 1’s. Ranges from 1-44. There are 47 missing values.
% RB (No. 1’s) Yield No 1’s divided by Russet Burbank’s yield No 1’s multiplied by 100. Ranges from 0-336. There are 47 missing values.
% No. 1’s Yield No 1’s divided by Yield Total multiplied by 100. Ranges from 0-93. There are 51 missing values.
$>$ 10 oz. Yield in cwt of potatoes above 10 oz. Sums of 10-20 oz (sometimes not provided) and over 20 oz. Ranges from 0-787. There are 143 missing values.
6-10 oz. Yield in cwt of potatoes between 6 and 10 oz. Ranges from 0-716. There are 144 missing values.
4-6 oz. Yield in cwt of potatoes between 4 and 6 oz. Ranges from 0-240. There are 143 missing values.
Yield No. 2’s + $>$ 20oz. Yield total minus Yield No. 1’s minus Yield Under 4 oz. minus Yield culls. Ranges from 0-429. There are 51 missing values.
Yield Under 4 oz. Yield in cwt of potatoes below 4 oz. Ranges from 0-807. There are 52 missing values.
Yield Culls Yield in cwt of discarded potatoes. Ranges from 0-178. There are 53 missing values.
Yield Over 20 oz. Yield in cwt of potatoes above 20 oz. There are 140 missing values.
Tuber/plant Total tuber count divided by the number of plants. There 276 missing values.
Average Tuber Size Measured in ounces. Ranges from 0-20. There are 175 missing values.
Length/Width Ratio Potato’s length divided by its width (computed from raw length and width, which are not separately included among the averages). Ranges from 0.52-3.5. There are 19 missing values.
Specific Gravity Potato’s air weight divided by the difference of its air and water weights (computed from raw air and water weights, which are not separately included among the averages). Ranges from 0-1.1. There are 18 missing values.
Fry Color Stem Measured in photovolts. Ranges from 14.4-72.9. There are 410 missing values.
Fry Color Bud Measured in photovolts. Ranges from 16.9-76.3. There are 410 missing values.
Hollow Heart Internal defect variable. Raw data is an expert-selected number from 1-10, multiplied by 10 to represent a percentage. Ranges from 0-90. There are 40 missing values.
Brown Center Internal defect variable. Raw data is an expert-selected number from 1-10, multiplied by 10 to represent a percentage. Ranges from 0-30. There are 40 missing values.
Black Spot Bruise Internal defect variable. Raw data is an expert-selected number from 1-10, multiplied by 10 to represent a percentage. Ranges from 0-87.5. There are 169 missing values.
Internal Brown Spot Internal defect variable. Raw data is an expert-selected number from 1-10, multiplied by 10 to represent a percentage. Ranges from 0-100. There are 40 missing values.
Vascular Discoloration Internal defect variable. Raw data is an expert-selected number from 1-10, multiplied by 10 to represent a percentage. Ranges from 0-100. There are 40 missing values.
Flower Color Plant characteristic. Ranges from 1-4 (white to pink). Variable scale. Should range from 1-5. There are 638 missing values.
Vine Size Plant characteristic. Ranges from 1.3-5.125. Variable scale. Should range from 1-5. There are 583 missing values.
Maturity Plant characteristic. Ranges from 0.1375-5.75. Variable scale. Should range from 1-5. There are 639 missing values.
Skin Color External characteristic. Ranges from 1.35-71.3. Variable scale. Should range from 1-5 (3 is yellow, 4 is brown). There are 290 missing values.
Russeting External characteristic. Ranges from 1-14.625. Variable scale. Should range from 1-5. There are 153 missing values.
Tuber Shape External characteristic. Ranges from 1-5 from spherical to long. There are 261 missing values.
Shape Uniformity External characteristic. Ranges from 1-5. There are 153 missing values.
Eye Depth External characteristic. Ranges from 1-5.5. Variable scale. Should range from 1-5. There are 119 missing values.
Greening External characteristic. Ranges from 1.625-5. Variable scale. Should range from 1-5. There are 153 missing values.
Growth Cracks External characteristic. Ranges from 1-5. There are 153 missing values.
Scab External characteristic. Ranges from 2.625-5. Variable scale. Should range from 1-5. There are 325 missing values.
Shatter Bruise External characteristic. Ranges from 2.16-20.83. Variable scale. Should range from 1-5. There are 176 missing values.
Bruise Data A string of number1 * number2 where number1 is the the diameter of the bruise and number2 is the depth of the bruise. This attribute is not included in the analysis as there is very little data for it.
[…]